[Brahms-l] Shift report 20040215 00:00-08:00

From: Apache <apache@bnl.gov>
Date: Sun Feb 15 2004 - 07:52:50 EST
------------------------------------------------------------
Date: Sun Feb 15 07:52:50 2004, Shift: 00:00-08:00
Supervisor: BHS, with: 
Log book#15 pages 58 - 58

Shift Summary:
------------------------------------------------------------
1) Good machine performance - the first store (00:00-03:25) was a bit low in initial intensity, so we voted dump together with PHENIX after 3hrs. The second one (04:30-still running) was very good.

2) Just before this shift, all magnets were flipped to A polarity and the FS was set to full field. D3 grumbled a bit and had to be tickled by the operator before it would hold current - reason unknown.
Runs: 10359 10360 10361 10362 10366 10367 10368
MRS: 45, 1050A
FS: 4, 1/1A
Trig2: 675k
Trig3: 1.17M (!) -> current MRS program is done! Will just keep going for the last hour of this store at least, then maybe get some more full field data?

3) No problems at all :-)
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sun Feb 15 07:53:05 2004

This archive was generated by hypermail 2.1.8 : Sun Feb 15 2004 - 07:53:24 EST